### v1

# Not All Samples Should Be Utilized Equally: Towards Understanding and Improving Dataset Distillation

Abstract
Dataset Distillation (DD) aims to synthesize a small dataset capable of performing comparably to the original dataset. Despite the success of numerous DD methods, theoretical exploration of this area remains unaddressed. In this paper, we take an initial step towards understanding various matching-based DD methods from the perspective of sample difficulty. We begin by empirically examining sample difficulty, measured by gradient norm, and observe that different matching-based methods roughly correspond to specific difficulty tendencies. We then extend the neural scaling laws of data pruning to DD to theoretically explain these matching-based methods. Our findings suggest that prioritizing the synthesis of easier samples from the original dataset can enhance the quality of distilled datasets, especially in low IPC (image-per-class) settings. Based on our empirical observations and theoretical analysis, we introduce the Sample Difficulty Correction (SDC) approach, designed to predominantly generate easier samples to achieve higher dataset quality. Our SDC can be seamlessly integrated into existing methods as a plugin with minimal code adjustments. Experimental results demonstrate that adding SDC generates higher-quality distilled datasets across 7 distillation methods and 6 datasets.
1 Introduction
In an era of data-centric AI, scaling laws [19] have shifted the focus to data quality. Under this scenario, dataset distillation (DD) [46, 39] has emerged as a solution for creating high-quality data summaries. Unlike data pruning methods [13, 7, 45, 1] that directly select data points from original datasets, DD methods are designed to generate novel data points through learning. The utility of DD methods has been witnessed in fields such as privacy protection [6, 29, 4, 11], continual learning [38, 15, 50, 32], neural architecture search [2, 34, 43], and federated learning [14, 28, 18].
Among the various DD techniques, matching-based methods, particularly gradient matching (GM) [52, 51, 25, 20] and trajectory matching (TM) [3, 8, 12, 16], have demonstrated outstanding performance. However, a gap remains between their theoretical understanding and empirical success. To offer a unified explanation of these methods, we aim to explore the following question:
Question 1: Is there a unified theory to explain existing matching-based DD methods?
Figure 1: We conducted experiments of GM-based methods on the FashionMNIST dataset and TM-based methods on the CIFAR-100 dataset. (a) Average gradient norms of network parameters for different GM-based methods are enhanced during distillation. The shade represents the gap before and after adding our SDC. (b) Test performance w/ and w/o our proposed SDC on different GM-based methods. Our SDC are incorporated to improve all matching-based methods. (c) Average gradient norms of network parameters for different TM-based methods are alleviated during distillation. (d) Test performance w/ and w/o our proposed SDC on different TM-based methods. Note that the average gradient norms are smoothed using the exponential moving average. Best viewed in color.
To address Question 1, we first empirically examine the differences between matching-based distillation methods. It is widely acknowledged that sample difficulty (Definition 1) is a crucial metric in data-centric AI that significantly affects model performance, as seen in dataset pruning [44, 31, 30, 42], active learning [40, 5, 49], and large language model prediction [9, 27, 26]. To track the differences between current distillation methods, we follow [37] and analyze sample difficulty using the GraDN metric (Definition 2). Surprisingly, we discover that the GraDN score is increased in GM-based methods (Figure 1(a)), while TM-based methods may reduce this metric (Figure 1(c)). These distinct trends indicate that the difficulty of samples utilized in GM-based methods is elevated (Figure 2(a)), whereas in TM-based methods, it is reduced (Figure 2(b)) during the distillation process.
Motivated by these empirical observations, we develop a theoretical explanation for current DD methods from the perspective of sample difficulty. Specifically, we draw upon the neural scaling law in the data pruning theory [42] to connect sample difficulty with distillation performance. As shown in Figure 4(c), our theory indicates that in matching-based DD methods, when the synthetic dataset is small—specifically, when the images-per-class (IPC) is low—the optimal strategy is to primarily focus on easier samples rather than harder ones to enhance data quality. Based on our theory, we further explain why TM-based methods usually outperform GM-based methods in real scenarios.
Beyond developing a theoretical framework, we take further steps to seek possible solutions for enhancing current approaches. This introduces another important research question:
Question 2: Is it empirically feasible to identify a loss function that surpasses the performance of the matching loss by controlling the difficulty of learned patterns during distillation?
To address Question 2, inspired by our empirical observations and theoretical analysis, we propose our novel Sample Difficulty Correction (SDC) method to enhance the synthetic dataset quality learned by current matching-based distillation methods. We achieve this by imposing the distillation method to focus more on learning from easy samples rather than hard samples. Specifically, we implicitly add a gradient norm regularizer to improve the synthetic dataset quality.
Our contributions are listed as follows:
  • We empirically investigate sample difficulty from the gradient norm perspective across various distillation methods, establishing a link between sample difficulty and synthetic dataset quality. We propose that GM-based methods tend to learn difficult samples during optimization, whereas TM-based methods do not exhibit a dominant preference for difficulty. This difference may explain the poorer performance of GM-based methods than TM-based methods in real-world applications.
  • We theoretically elucidate the mechanism of matching-based DD methods from the perspective of sample difficulty with statistical mechanics. Adapting the neural scaling law theory from data pruning [42] to distillation settings, we provide insights into how matching strategies evolve with the size of the synthetic dataset. Consequently, we propose that focusing on matching easy samples is a better strategy when the synthetic dataset is small.
  • We introduce Sample Difficulty Correction (SDC) to improve the quality of synthetic datasets in current matching-based DD methods. SDC can be incorporated into existing methods with minimal code modifications. Our method demonstrates superior generalization performance across 7 distillation methods (DC [52], DSA [51], DSAC [25], MTT [3], FTD [12], TESLA [8], and DATM [16]) and 6 datasets (MNIST [10], FashionMNIST [48], SVHN [36], CIFAR-10 [21], CIFAR-100 [21], and Tiny ImageNet [23]).
Figure 2: Comparison of matching-based dataset distillation methods from a sample difficulty perspective. (a) Gradient matching-based methods mainly utilize hard samples during synthesizing. (b) Trajectory matching-based methods do not explicitly take the difficulty of samples into consideration. (c) Our SDC encourages matching-based methods to prioritize the synthesis of easy samples.
2 Preliminaries and Related Work
Dataset distillation involves synthesizing a small, condensed dataset $\mathcal{D}_{\mathsf{syn}}$ that efficiently encapsulates the informational essence of a larger, authentic dataset $\mathcal{D}_{\mathsf{real}}$.
Gradient Matching (GM) based methods are pivotal in achieving distillation by ensuring the alignment of training gradients between surrogate models trained on both the original dataset $\mathcal{D}_{\mathsf{real}}$ and the synthesized dataset $\mathcal{D}_{\mathsf{syn}}$. This method is first introduced by DC [52]. Let $\theta_{t}$ represent the network parameters sampled from distribution $P_{\theta}$ at step $t$, and $C$ symbolizes the categories within $\mathcal{D}_{\mathsf{real}}$. The cross-entropy loss, denoted by $\mathcal{L}$, is employed to assess the matching loss by comparing the gradient alignment over a time horizon of $T$ steps. The formal optimization objective of DC is:
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}{\operatorname{\arg\min}}\underset{\theta_{0}\sim P_{\theta},c\sim C}{\mathbb{E}}\left[\sum_{t=0}^{T}\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right)\right],$ (1)
where $\mathbf{D}$ measures the cumulative distances (e.g., cosine/$L_{2}$ distance in DC) between the gradients of weights corresponding to each category output. The parameter updates for $\theta$ are executed in an inner loop via gradient descent, with a specified learning rate $\eta$:
$\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\left(\theta_{t}\right).$ (2)
Building upon this, DSA [51] enhances DC by implementing consistent image augmentations on both $\mathcal{D}_{\mathsf{real}}$ and $\mathcal{D}_{\mathsf{syn}}$ throughout the optimization process. Moreover, DCC [25] refines the gradient matching objective by incorporating class contrastive signals at each gradient matching step, which results in enhanced stability and performance. Combining DSA and DCC, DSAC [25] further introduces improvements by synergizing these techniques. The revised optimization objective for DCC and DSAC is formulated as:
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}{\operatorname{\arg\min}}\underset{\theta_{0}\sim P_{\theta}}{\mathbb{E}}\left[\sum_{t=0}^{T}\mathbf{D}\left(\mathbb{E}_{c\in C}\left[\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}(\theta_{t})\right],\mathbb{E}_{c\in C}\left[\nabla_{\theta}\mathcal{L}_{\mathcal{D}^{c}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}(\theta_{t})\right]\right)\right].$ (3)
Trajectory matching (TM) based approaches aim to match the training trajectories of surrogate models by optimizing over both the real dataset $\mathcal{D}_{\mathsf{real}}$ and the synthesized dataset $\mathcal{D}_{\mathsf{syn}}$. TM-based methods were initially proposed in MTT [3]. Let term $\tau^{\mathcal{D}_{\mathsf{real}}}$ denote the expert training trajectories, represented as a sequential array of parameters $\{\theta_{t}^{\mathcal{D_{\mathsf{real}}}}\}_{t=0}^{T}$, obtained from training a network on the real dataset $\mathcal{D}_{\mathsf{real}}$. In parallel, ${\theta}_{t}^{\mathcal{D}_{\mathsf{syn}}}$ refers to the parameter set of the network trained on $\mathcal{D}_{\mathsf{syn}}$ at step $t$. In each iteration, parameters $\theta_{t}^{\mathcal{D}_{\mathsf{real}}}$ and $\theta_{t+M}^{\mathcal{D}_{\mathsf{real}}}$ are randomly selected from the expert trajectory pool $\{\tau^{\mathcal{D}_{\mathsf{real}}}\}$, serving as the initial and target parameters for trajectory alignment, where $M$ is a predetermined hyperparameter. TM-based methods enhance the synthetic dataset $\mathcal{D}_{\mathsf{syn}}$ by minimizing the loss defined as:
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}{\arg\min}\underset{\theta_{0}\sim P_{\theta}}{\mathbb{E}}\left[\sum_{t=0}^{T-M}\frac{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t+N}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\right)}{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}\right)}\right],$ (4)
where $\mathbf{D}$ is a distance metric (e.g., $L_{2}$ distance in MTT) and $N<<M$ is a predefined hyperparameter. ${\theta}_{t+N}^{\mathcal{D}_{\mathsf{syn}}}$ is derived through an inner optimization using the cross-entropy loss $\mathcal{L}$ with the learning rate $\eta$:
${\theta}_{t+i+1}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\leftarrow{\theta}_{t+i}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}-\eta\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}({\theta}_{t+i}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}),\text{ where }{\theta}_{t}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}:=\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}.$ (5)
Along this line, TESLA [8] utilizes linear algebraic manipulations and soft labels to increase compression efficiency, FTD [12] aims to seek a flat trajectory to avoid accumulated trajectory error, and DATM [16] considers matching only necessary parts of trajectory with difficulty alignment.
3 Method
Figure 3: The statistical relationship between sample difficulty $\chi(x,y;\Theta_{t})$, gradient norm $\text{GraDN}(x,y;\Theta_{t})$, and average validation loss for each sample $(x,y)$ on a series of models with $\theta_{t}\in\Theta_{t}$. We observe a significant positive correlation between sample difficulty and both the gradient norm and the loss. Experiments were conducted using ResNet-18 on CIFAR-10 and ResNet-34 on CIFAR-100. Each score was evaluated across 20 pretrained models. For CIFAR-100, 10 categories were randomly selected for visualization. The relationships depicted are: (a) sample difficulty vs. gradient norm on CIFAR-10, (b) sample difficulty vs. loss on CIFAR-10, (c) sample difficulty vs. gradient norm on CIFAR-100, and (d) sample difficulty vs. loss on CIFAR-100.
3.1 A Closer Look at Sample Difficulty
In this subsection, we aim to intuitively understand dataset distillation through the concept of sample difficulty (Definition 1), which is pivotal in data-centric AI [44, 31, 30, 5, 49, 9, 26]. We begin by empirically observing the evolution of sample difficulty during the distillation process. Firstly, we introduce the commonly used definition of sample difficulty, namely the GraDN score (Definition 2), and validate the reliability of this metric. Furthermore, we track the GraDN score across current dataset distillation methods to delve deeper into their underlying mechanisms.
Definition 1 (Sample Difficulty [33]).
Given a training pair $(x,y)$ and a series of pretrained models at training time $t$, the sample difficulty, denoted $\chi(x,y;\Theta_{t})$, is defined as the expected probability of $(x,y)$ being misclassified by an ensemble of models $\theta_{t}\in\Theta_{t}$. Formally, it is presented as:
$\chi(x,y;\Theta_{t})=\mathbb{E}_{\theta_{t}\in\Theta_{t}}\left[\mathbbm{1}\left(y\neq\theta_{t}(x)\right)\right],$ (6)
where $\mathbbm{1}\left(z\right)$ is an indicator function that equals 1 if the boolean input $z$ is true, and 0 otherwise. In this case, the indicator function equals to 1 if the sample $(x,y)$ is misclassified by the model with parameters $\theta_{t}$, and 0 otherwise.
Definition 2 (GraDN Score [37]).
Consider a training pair $(x,y)$, with $\mathcal{L}$ representing the loss function. At time $t$, the GraDN score for $(x,y)$ is calculated as the average gradient norm of the loss $\mathcal{L}$ across a diverse ensemble of models with parameters $\theta_{t}\in\Theta_{t}$:
$\operatorname{GraDN}(x,y;\Theta_{t})=\mathbb{E}_{\theta_{t}\in\Theta_{t}}\left[\|\nabla_{\theta}\mathcal{L}(x,y;\theta_{t})\|_{2}\right],$ (7)
where $\nabla_{\theta}\mathcal{L}(x,y;\theta_{t})$ denotes the gradient of loss $\mathcal{L}$ on sample $(x,y)$ w.r.t. the model parameters $\theta_{t}$, and $\|\cdot\|_{2}$ denotes $L_{2}$ norm.
According to [33], the difficulty of each sample can be assessed by the misclassification ratio across a series of pretrained models (Definition 1). Additionally, from an optimization perspective, it can be represented by the gradient norm of the loss on a series of pretrained models for this sample (Definition 2). In our study, we adopt Definition 2 to evaluate sample difficulty as interpreted by various matching methods.
Empirical verification of the relationship between sample difficulty and gradient norm. We conducted experiments to verify the reliability of the GraDN score in classifying the CIFAR-10 and CIFAR-100 datasets by training a set of models. As depicted in Figure 3, the GraDN score shows a clear positive correlation with the sample difficulty. For easier samples, GraDN scores are generally lower, exerting minimal impact on the network’s gradient flow. Conversely, for harder samples, higher GraDN scores indicate a significant impact on the optimization directions of the models. We show detailed results of the relationships between these metrics in Appendix A.
Exploring sample difficulty across different distillation methods. Beyond sample difficulty under classification scenarios, we now extend our observations to matching-based distillation methods. Specifically, we examined the average gradient norm of the training cross-entropy loss across network parameters during the distillation process. As shown in Figure 1(a)(c), we found that the average gradient norm (corresponding to the GraDN score) tends to increase in GM-based methods (signifying harder samples), whereas it decreases in TM-based methods (indicating easier samples). This unexpected phenomenon motivates us to further theoretically explore matching-based distillation methods from the perspective of sample difficulty.
3.2 An Analytical Theory for Explaining Matching-based Dataset Distillation
In Section 3.1, we empirically observed distinct trends in sample difficulty across various dataset distillation methods. Here, we propose an analytical theory based on the neural scaling law to formally analyze sample difficulty in matching-based methods. We extend the theory of data pruning presented by [42] and validate its applicability within the context of DD using an expert-student perceptron model. Unlike data pruning, where the pruned dataset is directly selected from the original dataset, DD involves synthesizing a small, new, unseen dataset.
Figure 4: An analytic theory of dataset distillation. (a) Test error $\varepsilon$ as a function of the synthetic samples per parameter $\alpha_{\mathsf{syn}}$ and fraction of data synthesized $f$ in a perfect expert setting ($\gamma=0$). (b) We show the difference in test accuracy in learning synthetic dataset by learning from hard vs easy samples, revealing the change in distillation strategy. (c) Our theory suggests that when we synthesize small dataset (small $\alpha_{\mathsf{syn}}$), the better distillation strategy is to utilize the easy samples.
We start our analysis with tools from statistical mechanics [35]. Let us consider a classification problem in dataset $\mathcal{D}^{\mathsf{real}}$ containing $d_{\mathsf{real}}$ samples $\{x_{i},y_{i}\}_{i=1,\ldots,d_{\mathsf{real}}}$, where $x_{i}\in\mathbb{R}^{d}\sim\mathcal{N}(0,I_{d})$ are i.i.d. zero-mean, unit variance Gaussian inputs, and $y_{i}=\text{sign}(\theta^{\mathcal{D}_{\mathsf{real}}^{\top}}x_{i})\in\{-1,+1\}$ are labels generated by an expert perceptron $\theta^{\mathcal{D}_{\mathsf{real}}}\in\mathbb{R}^{d}$. Our analysis is within the high-dimensional statistics limit, where $d,d_{\mathsf{real}}\to\infty$ while maintaining the ratio of total training samples to parameters $\alpha_{\mathsf{tot}}=d_{\mathsf{real}}/d$ at $O(1)$. The general distillation algorithm proceeds as follows:
  1. 1.
    Train a student perceptron on $\mathcal{D}_{\mathsf{real}}$ for a few epochs to obtain weights $\theta^{\mathsf{probe}}$. The gap between can be measured by the angle $\gamma$ between the probe student $\theta^{\mathsf{probe}}$ and the expert $\theta^{\mathsf{real}}$. If $\theta^{\mathsf{probe}}\approx\theta^{\mathsf{real}}$, we denote the $\theta^{\mathsf{probe}}$ as a perfect probe ($\gamma=0$). Otherwise, in imperfect probe cases, $\gamma\neq 0$.
  2. 2.
    Compute the margin $m_{i}=\theta^{\mathsf{probe}^{\top}}(y_{i}x_{i})$ for each training example, categorizing large (small) margins as easy (hard) samples.
  3. 3.
    Generate a synthetic dataset $\mathcal{D}_{\mathsf{syn}}$ of size $d_{\mathsf{syn}}=fd_{\mathsf{real}}$, by learning from the hardest samples from $\mathcal{D}_{\mathsf{real}}$ for few steps. The ratio of total samples of the synthetic dataset to parameters is $\alpha_{\mathsf{syn}}=d_{\mathsf{syn}}/d$.
  4. 4.
    Train a new perceptron $\theta^{\mathcal{D}_{\mathsf{syn}}}$ on $\mathcal{D}_{\mathsf{syn}}$, aiming to classify training data with the maximal margin $\kappa=\min_{i}\theta^{\mathcal{D}_{\mathsf{syn}}^{\top}}y_{i}x_{i}$.
We analyze the test error $\varepsilon$ of this final perceptron $\theta^{\mathcal{D}_{\mathsf{syn}}}$ as a function of $\alpha_{\mathsf{syn}}$, $f$, and the angle $\gamma$ between the probe student $\theta^{\mathsf{probe}}$ and the expert $\theta^{\mathcal{D}_{\mathsf{real}}}$. Similar to [42], we define $\theta^{\mathsf{probe}}$ as a random Gaussian vector conditioned to have angle $\gamma$ with the expert. Under this scenario, we can derive an asymptotically exact analytic expression for $\varepsilon(\alpha_{\mathsf{syn}},f,\gamma)$ (see Appendix D for details):
$\varepsilon=\frac{\cos^{-1}(R)}{\pi},\quad\text{where }R=\frac{\theta^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{\top}}\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}}{\|\theta^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}\|_{2}\|\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\|_{2}}$ (8)
Likewise in [42], we can solve $R$ with saddle point equations in Appendix D.1, enabling direct predictions of the test error $\varepsilon$ according to Eq (8).
Verification of the neural scaling law in dataset distillation. We first evaluated the correctness of our theory in the perfect expert-student setting ($\gamma=0$). As shown in Figure 4(a), we observed an excellent match between our analytic theory (solid curves) and numerical simulations (dots) of perceptron learning at parameters $d=200$ in dataset distillation. We also verify our theory in imperfect probe settings when $\gamma\neq 0$, as shown in Appendix D.2.
The neural scaling law for dataset distillation. We further investigate the relationship between the distillation ratio ($\alpha_{\mathsf{syn}}$), the fraction of data synthesized ($f$), and the final accuracy of $\theta^{\mathcal{D}_{\mathsf{syn}}}$ under various distillation strategies, such as synthesizing $\mathcal{D}_{\mathsf{syn}}$ from hard or easy samples. Similar to data pruning, when $f=1$ (no distillation), the test error follows the classical perceptron learning power-law scaling, $\varepsilon\propto\alpha_{\mathsf{syn}}^{-1}$. In other cases, our findings reveal that for smaller $\alpha_{\mathsf{syn}}$ (smaller synthetic datasets), learning from the hard samples results in poorer performance than no distillation. Conversely, for larger $\alpha_{\mathsf{syn}}$, focusing on the hard samples yields substantially better outcomes than no distillation. We find that in limited data regimes, matching the easy samples, which have the largest margins, offers a more effective distillation strategy. This finding highlights that in most cases of DD (where $d_{\mathsf{syn}}<<d_{\mathsf{real}}$), it is crucial for the model to first learn from the basic characteristics in $\mathcal{D}_{\mathsf{real}}$; hence, prioritizing easy samples facilitates reaching a moderate error level more swiftly.
Understanding GM-based and TM-based methods with the neural scaling law. As depicted in Figure 1(a), we observe that GM-based methods typically incorporate the hard samples within the synthetic dataset. This trend is due to the GM-based matching loss in Eq.( 1), which predominantly penalizes samples with large gradients (hard samples) (shown in Figure 2(b)). However, in most DD settings, the size of synthetic dataset $\mathcal{D}_{\mathsf{syn}}$ is usually small. Therefore, according to our theory, we should mainly focus on synthesizing the dataset by matching easy samples to achieve higher dataset quality. In contrast, the simplicity of the synthetic samples generated by TM-based methods, as shown in Figure 1(a), is not directly concerned through distillation. From Eq.( 4), it is evident that TM-based methods prioritize parameter alignment, thus penalizing the matching term without explicitly targeting sample difficulty (shown in Figure 2(c)). This approach results in a synthetic dataset that may be generated by learning samples of randomly vary in difficulty. We can provide a explanation that TM-based methods generalize well in real scenarios than GM-based methods because of they do not explicitly focus on synthesizing by matching hard samples.
3.3 Matching with Sample Difficulty Correction
Based on our theoretical analysis of matching-based dataset distillation, we propose a novel method to enhance existing techniques for synthesizing higher-quality distilled datasets. Although TM-based methods have achieved relative success on current benchmark datasets, they do not explicitly consider sample difficulty, which could ensure higher synthetic dataset quality.
A direct approach to impose constraints on sample difficulty is to calculate the gradient norm for each sample as a metric to determine its utility. Let us consider the case of GM-based methods. At step $t$, a batch of real samples $\mathcal{B}_{\mathsf{real}}^{c}\sim\mathcal{D}_{\mathsf{real}}^{c}$ of class $c\in C$ is to be matched with the gradients of a synthetic batch $\mathcal{B}_{\mathsf{syn}}^{c}\sim\mathcal{D}_{\mathsf{syn}}^{c}$. To decide whether to utilize each sample in $\mathcal{B}_{\mathsf{real}}^{c}$, it is natural to compute the gradient norm of each sample and utilize those with a score smaller than a predefined threshold $\tau$. Specifically, a sample $(x,y)$ is utilized if $\|\nabla_{\theta}\mathcal{L}(x,y;\theta_{t})\|_{2}\leq\tau$. Consequently, the modified loss for matching only easy samples is:
~$\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right],\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right],$ (9)
where ~$\tilde{\mathcal{B}}_{\mathsf{real}}^{c}=\{(x,y)|(x,y)\in\mathcal{B}_{\mathsf{real}}^{c},\|\nabla_{\theta}\mathcal{L}(x,y;\theta_{t})\|_{2}\leq\tau\}$ denotes the modified batch with only easy samples, and ~$\tilde{\mathcal{B}}_{\mathsf{syn}}^{c}$ denote a sampled batch from $\mathcal{D}_{\mathsf{syn}}^{c}$ with the same size as ~$\tilde{\mathcal{B}}_{\mathsf{real}}^{c}$. The corresponding matching loss should be:
$\tilde{L}(\theta_{t})=\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\tilde{\mathcal{B}}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right),$ (10)
However, the computational cost of constructing reduced easy sample batch ~$\tilde{\mathcal{B}}_{\mathsf{real}}^{c}$ from $\mathcal{B}_{\mathsf{real}}^{c}$ is unrealistic in real-world scenarios because it requires calculating the gradient norm for each sample independently, resulting in a tenfold or greater increase in time. Besides, determining the difficulty threshold $\tau$ is also ad-hoc and challenging for each sample. Therefore, we take an alternative approach, i.e., we consider adding the overall sample difficulty of the whole batch $\mathcal{B}_{\mathsf{syn}}^{c}$ as an implicit regularization term in the matching loss function. Our proposed methods, named Sample Difficulty Correction (SDC), can be incorporated into current matching methods with minimal adjustment of code implementation. Specifically, for a single-step GM, we have the following modified loss:
$L_{\lambda}(\theta_{t})=\underbrace{\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right)}_{\text{Gradient Matching Loss}}+\underbrace{\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\right\|_{2}}_{\text{Gradient Norm Regularization}}$ (11)
For TM-based methods that do not explicitly focus on sample difficulty during distillation, we compute the average gradient norm of the whole dataset $\mathcal{D}_{\mathsf{syn}}$ during the optimization of the student network $\theta_{t}^{\mathcal{D}_{\mathsf{syn}}}$ w.r.t. the training loss as the regularization term. Specifically, we have:
$L_{\lambda}(\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}_{t})=\underbrace{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t+N}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\right)/{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}\right)}}_{\text{Trajectory Matching Loss}}+\underbrace{\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\right\|_{2}}_{\text{Gradient Norm Regularization}}$ (12)
By adding the gradient norm regularization, we can implicitly enforce current matching-based methods to mainly concentrate on synthesizing easy samples to achieve better synthetic data quality. We provide the detailed algorithm pseudocodes for GM- and TM-based methods in Appendix B.3.
Table 1: Comparison of test accuracy (%) results of GM-based dataset distillation methods w/ and w/o SDC on MNIST, FashionMNIST, and SVHN datasets.
Dataset MNIST FashionMNIST SVHN
IPC 1 10 50 1 10 50 1 10 50
Ratio (%) 0.02 0.2 1 0.2 2 10 0.2 2 10
Random 64.9±3.5 95.1±0.9 97.9±0.2 51.4±3.8 73.8±0.7 82.5±0.7 14.6±1.6 35.1±4.1 70.9±0.9
Herding 89.2±1.6 93.7±0.3 94.8±0.2 67.0±1.9 71.1±0.7 71.9±0.8 20.9±1.3 50.5±3.3 72.6±0.8
Forgetting 35.5±5.6 68.1±3.3 88.2±1.2 42.0±5.5 53.9±2.0 55.0±1.1 12.1±1.7 16.8±1.2 27.2±1.5
DC 91.8±0.4 97.4±0.2 98.5±0.1 70.3±0.7 82.1±0.3 83.6±0.2 31.1±1.3 75.3±0.6 82.1±0.2
+SDC 92.0±0.4 97.5±0.1 98.9±0.1 70.7±0.5 82.4±0.3 84.7±0.2 31.4±1.2 76.0±0.5 82.3±0.3
DSA 88.9±0.8 97.2±0.1 99.1±0.1 70.1±0.4 84.7±0.2 88.7±0.2 29.4±1.0 79.2±0.4 84.3±0.4
+SDC 89.2±0.4 97.3±0.1 99.2±0.4 70.5±0.5 84.8±0.2 88.9±0.1 30.6±1.0 79.4±0.4 85.3±0.4
DSAC 89.2±0.7 97.7±0.1 98.8±0.1 71.8±0.7 84.9±0.2 88.5±0.2 47.5±1.8 80.1±0.5 87.3±0.2
+SDC 89.7±0.7 97.8±0.1 98.9±0.1 72.2±0.6 85.1±0.1 88.7±0.1 48.1±1.6 80.4±0.3 87.4±0.2
Whole Dataset 99.6±0.0 93.5±0.1 95.4±0.1
4 Experiments
4.1 Basic Settings
Datasets and baselines. For GM-based methods, we followed previous works to conduct experiments on MNIST [10], FashionMNIST [48], SVHN [36] datasets. We utilized current GM-based methods, including DC [52], DSA [51], and DSAC [25] as baselines. For TM-based methods, we followed the recent papers to use CIFAR-10, CIFAR-100 [21], and Tiny ImageNet [23] datasets. We performed experiments on current baselines including MTT [3], FTD [12], TESLA [8], and DATM [16]. We added our Sample Difficulty Correction (SDC) for all these baseline methods. To ensure a fair comparison, we employed identical hyperparameters for GM-based and TM-based methods with and without SDC while keeping all other variables constant, such as model architecture and augmentations. As per convention, for TM-based methods, we used max test accuracy, while for GM-based methods, we utilized the test accuracy from the last iteration. We also compared our methods with classical data pruning algorithms including Random, Herding [47], and Forgetting [45]. All hyperparameters are detailed in Appendix B.1.
Neural networks for distillation. Without specific mention, we used ConvNet as default to conduct experiments. Consistent with other previous methods, we used 3-layer ConvNet for CIFAR-10, CIFAR-100, MNIST, SVHN, and FashionMNIST, and 4-layer ConvNet for Tiny ImageNet.
4.2 Main Results
GM-based methods on MNIST, FashionMNIST, and SVHN. As presented in Table 1, we report the results of three GM-based methods applied to MNIST, FashionMNIST, and SVHN datasets. Each method was evaluated with IPC (images-per-class) values of 1, 10, and 50. Notably, adding SDC improves the test accuracy of baseline methods across all datasets and IPC values, demonstrating the effectiveness of our approach. Notably, adding SDC to the original method improved the test accuracy of DSA by 1.2% on the SVHN dataset with IPC = 1, and by 1% with IPC = 50. For DC on the FashionMNIST dataset with IPC = 50, the test accuracy was increased by 1.1% with SDC. All hyperparameters are detailed in Table 4.
TM-based methods on CIFAR-10/100 and Tiny ImageNet. As shown in Table 2, we present the results of four TM-based methods trained on CIFAR-10, CIFAR-100 and Tiny ImageNet. By incorporating the average gradient norm as a regularization term during matching with SDC, the resulting test accuracy was generally improved. Notably, employing SDC improved the test accuracy of FTD on CIFAR-10 by 1.2% with IPC = 10 and 1.1% with IPC = 50, and enhanced the test accuracy of DATM on Tiny ImageNet by 0.6%. For FTD, we used EMA (exponential moving average) just as in the original method[12]. All hyperparameters are detailed in Tables 567, and 8.
Table 2: Comparison of test accuracy (%) results of TM-based dataset distillation methods w/ and w/o SDC on CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
Dataset CIFAR-10 CIFAR-100 Tiny ImageNet
IPC 1 10 50 1 10 50 1 10 50
Ratio (%) 0.02 0.2 1 0.2 2 10 0.2 2 10
Random 14.4±2.0 26.0±1.2 43.4±1.0 4.2±0.3 14.6±0.5 30.0±0.4 1.4±0.1 5.0±0.2 15.0±0.4
Herding 21.5±1.2 31.6±0.7 40.4±0.6 8.4±0.3 17.3±0.3 33.7±0.5 2.8±0.2 6.3±0.2 16.7±0.3
Forgetting 13.5±1.2 23.3±1.0 23.3±1.1 4.5±0.2 15.1±0.3 30.5±0.3 1.6±0.1 5.1±0.2 15.0±0.3
MTT 45.8±0.3 64.7±0.5 71.5±0.5 23.9±1.0 38.7±0.4 47.3±0.1 8.3±0.4 20.6±0.2 28.0±0.3
+SDC 46.2±0.7 65.3±0.3 71.8±0.5 24.3±0.3 38.8±0.3 47.3±0.2 8.5±0.2 20.7±0.2 28.0±0.2
FTD 46.7±0.7 65.2±0.5 72.2±0.1 25.1±0.4 42.5±0.1 50.3±0.3 10.9±0.1 21.8±0.3 -
+SDC 47.2±0.7 66.4±0.4 73.3±0.4 25.4±0.3 42.6±0.1 50.5±0.3 11.2±0.1 22.2±0.2 -
TESLA 47.4±0.3 65.0±0.7 71.4±0.5 23.9±0.3 35.8±0.7 44.9±0.4 - - -
+SDC 47.9±0.7 65.3±0.4 71.8±0.2 24.2±0.2 35.9±0.2 45.0±0.4 - - -
DATM 46.1±0.5 66.4±0.6 75.9±0.3 27.7±0.3 47.6±0.2 52.1±0.1 17.1±0.3 30.1±0.3 39.7±0.1
+SDC 46.4±0.4 66.6±0.4 76.1±0.2 28.0±0.2 47.8±0.2 52.5±0.2 17.4±0.2 30.7±0.2 39.9±0.2
Whole Dataset 84.8±0.1 56.2±0.3 37.6±0.4
Generalization performance to other architectures. We evaluated the generalizability of synthetic datasets generated through distillation. Specifically, we used DSAC and DATM, which are current SOTA methods in GM-based and TM-based distillation, respectively. After distillation, the synthetic datasets were assessed using various neural networks, including ResNet-18 [17], VGG-11 [41], AlexNet [22], LeNet [24] and MLP. As shown in Table 3, even though our synthetic datasets were distilled using ConvNet, it generalizes well across most networks. Notably, for the experiment of DATM on CIFAR-10 with IPC = 1, employing SDC resulted in an accuracy improvement of 4.61% when using AlexNet. Employing SDC to DSAC led to an accuracy improvement of 0.9% on SVHN with IPC = 10 when using MLP. Additional results can be found in Appendix C.1.
Table 3: Cross-architecture evaluation was conducted on the distilled dataset with (a) IPC = 1 for the TM-based method (DATM) and (b) IPC = 10 for the GM-based method (DSAC), both w/ and w/o SDC. Adding SDC improves performance on unseen networks compared to current SOTA methods.
(a)  Dataset Method ResNet-18 VGG-11 AlexNet LeNet MLP CIFAR-10 DATM 29.62 25.12 19.38 23.41 23.08 +SDC 31.29 25.99 23.99 23.65 22.90 CIFAR-100 DATM 11.52 8.74 1.95 6.71 6.47 +SDC 12.10 8.78 3.73 6.84 6.51 Tiny ImageNet DATM 4.36 5.93 4.33 2.42 2.29 +SDC 4.74 6.45 4.34 2.79 2.32
(b)  Dataset Method ResNet-18 VGG-11 AlexNet LeNet MLP MNIST DSAC 97.44 96.88 95.30 95.31 90.62 +SDC 97.65 97.13 95.73 95.67 90.93 FashionMNIST DSAC 82.17 82.59 80.73 79.82 80.09 +SDC 82.87 82.73 81.15 79.96 80.36 SVHN DSAC 70.59 76.44 49.66 55.98 39.11 +SDC 71.03 76.63 49.78 56.27 40.03
4.3 Further Discussions
Discussion of SDC coefficient $\lambda$. The selection of the regularization coefficient $\lambda$ is pivotal for the quality of the distilled dataset. Our theory suggests that a larger $\lambda$ typically produces better synthetic datasets for smaller IPC values. Ideally, for low IPC settings, it is better to employ a large $\lambda$ to strongly penalize sample difficulty, whereas, for high IPC settings, the required $\lambda$ can be small or even close to zero in extreme cases. For simplicity and to maintain consistency across different datasets and baseline methods, we have set $\lambda=0.002$ as the default value in most of our experiments. As demonstrated in Figure 6, this choice of $\lambda$ aligns with the IPC values. Results for FTD and TESLA are based on CIFAR-10, results for DSA are based on SVHN, and results for DSAC are based on MNIST. Additionally, we further show that the choice of $\lambda$ is not sensitive in Appendix C.3.
Adaptive sample difficulty correction by adaptively increasing $\lambda$ during distillation. While our SDC seeks simplicity in regularization, DATM [16] claims that the matching difficulty is increased through optimization. Inspired by their observation, we implemented a strategy where $\lambda$ increases progressively throughout the matching phases. This method is designed to incrementally adjust the focus from easier to more complex patterns. Inspired by their observation, we applied an Adaptive Sample Difficulty Correction (ASDC) strategy in our experiments with a TM-based method on the CIFAR-100 with IPC = 1 and with a GM-based method on the FashionMNIST with IPC = 1. The $\lambda$ of DATM was initialized to 0.02 and logarithmically increased to 0.08 over 10,000 iterations and DSAC was initialized to 0.002 and logarithmically increased to 0.008 over 10,000 steps. For DATM, we use max test accuracy, while for DSAC, we use test accuracy. Experimental results of ASDC validate its potential to significantly enhance learning by finetuning regularization according to the complexity of the learned patterns. Figure 6 illustrates that ASDC further improves our method within SOTA matching methods. Additional results are provided in Appendix C.2.
Figure 5: Optimal $\lambda$ values for various matching-based distillation methods with SDC, performed on datasets with different IPC values.
Figure 6: Flexibly adjusting sample difficulty correction with an adaptive increase in $\lambda$ results in higher accuracy compared to standard SDC and baseline methods. We present the results of (a) DATM and (b) DSAC.
5 Conclusion
In this study, we empirically examine the matching-based dataset distillation method in relation to sample difficulty, observing clear trends as measured by gradient norm. Additionally, we adapt a neural scaling law from data pruning to theoretically explain dataset distillation. Our theoretical analysis suggests that for small synthetic datasets, the optimal approach is to generate data using easier samples from the original dataset rather than harder ones. To facilitate this, we propose a simplicity-centric regularization method, termed Sample Difficulty Correction (SDC), aimed at improving synthetic data quality by predominantly utilizing easier samples in the data generation process. This method can be easily incorporated to existing matching-based methods, and can be implemented with a few lines of code. Experimental results underscore the importance of proper regularization within the optimization process for dataset distillation. We anticipate that this work will deepen the theoretical understanding of dataset distillation.
Appendix A More Result on the Relationships between Sample Difficulty, Gardient Norm, and Loss
In this section, we present further findings on the relationships between sample difficulty $\chi(x,y;\Theta_{t})$, gradient norm $\text{GraDN}(x,y;\Theta_{t})$, and the average validation loss for each sample $(x,y)$ across a range of models characterized by $\theta_{t}\in\Theta_{t}$. The experiments were conducted using ResNet-18 on CIFAR-10 and ResNet-34 on CIFAR-100. Each metric was evaluated across 20 pretrained models. We randomly selected 1000 samples for each category in CIFAR-10, and 100 samples for each category in CIFAR-100. For CIFAR-100, 10 categories for visualization were randomly selected for visualization purposes. As shown in Figure 7, Figure 8, Figure 9, Figure 10, Figure 11, Figure 12, it reveals a significant positive correlation between sample difficulty, gradient norm and loss.
Figure 7: The statistical relationship between sample difficulty $\chi(x,y;\Theta_{t})$, gradient norm $\text{GraDN}(x,y;\Theta_{t})$ for each sample $(x,y)$ on a series of ResNet-18 models with parameters $\theta_{t}\in\Theta_{t}$ on CIFAR-10. 1000 samples were randomly selected for each category.
Figure 8: The statistical relationship between gradient norm $\text{GraDN}(x,y;\Theta_{t})$ and average validation loss for each sample $(x,y)$ on a series of ResNet-18 models with parameters $\theta_{t}\in\Theta_{t}$ on CIFAR-10. 1000 samples were randomly selected for each category.
Figure 9: The statistical relationship between sample difficulty $\chi(x,y;\Theta_{t})$ and gradient norm $\text{GraDN}(x,y;\Theta_{t})$ for each sample $(x,y)$ on a series of ResNet-18 models with parameters $\theta_{t}\in\Theta_{t}$ on CIFAR-10. 1000 samples were randomly selected for each category.
Figure 10: The statistical relationship between sample difficulty $\chi(x,y;\Theta_{t})$, gradient norm $\text{GraDN}(x,y;\Theta_{t})$ for each sample $(x,y)$ on a series of ResNet-34 models with parameters $\theta_{t}\in\Theta_{t}$ on CIFAR-100. 100 samples were randomly selected for each category.
Figure 11: The statistical relationship between gradient norm $\text{GraDN}(x,y;\Theta_{t})$ and average validation loss for each sample $(x,y)$ on a series of ResNet-34 models with parameters $\theta_{t}\in\Theta_{t}$ on CIFAR-100. 100 samples were randomly selected for each category.
Figure 12: The statistical relationship between sample difficulty $\chi(x,y;\Theta_{t})$ and gradient norm $\text{GraDN}(x,y;\Theta_{t})$ for each sample $(x,y)$ on a series of ResNet-34 models with parameters $\theta_{t}\in\Theta_{t}$ on CIFAR-100. 100 samples were randomly selected for each category.
Appendix B More Details of Experiments
B.1 Parameter Tables
B.1.1 GM-based Methods
Regarding the GM-based methods, Table 4 provides the corresponding $\lambda$ values after applying SDC. All results are obtained from a single experiment, and evaluated 20 times. Baseline results are obtained using identical configurations with the original methods’ implementations (please refer to DC and DSA [^1] https://github.com/VICO-UoE/DatasetCondensation, and DCC [^2] https://github.com/Saehyung-Lee/DCC). Experiments with our SDC share consistent hyperparameters with the corresponding baselines.
Table 4: The $\lambda$ used for GM-based methods
Dataset MNIST FashionMNIST SVHN
IPC 1 10 50 1 10 50 1 10 50
DC 0.001 0.0005 0.001 0.0002 0.001 0.01 0.001 0.0005 0.0002
DSA 0.001 0.00002 0.001 0.0002 0.002 0.001 0.005 0.0005 0.01
DSAC 0.001 0.02 0.02 0.002 0.02 0.02 0.005 0.02 0.02
B.1.2 TM-based Methods
The hyperparameters used in our TM-based methods differ slightly from the original methods (see original implementations of MTT [^3] https://github.com/GeorgeCazenavette/mtt-distillation, DATM [^4] https://github.com/NUS-HPC-AI-Lab/DATM, TESLA [^5] https://github.com/justincui03/tesla, and FTD [^6] https://github.com/AngusDujw/FTD-distillation), particularly in terms of synthesis steps, number of evaluations, and evaluation interval. Our baseline results used the settings in Table 5, Table 6, Table 7 and Table 8. The experiments of applying SDC were conducted in the same setting as in the baselines. In Table 6, and Table 7, we report the optimal hyperparameters using the ConvNetD3 network. All combinations in Table 7 and Table 8 used the ZCA.
Table 5: Optimal hyperparameters for MTT. A synthesis batch size of “-” means that we used the full support set at each synthesis step.
Dataset Model IPC ZCA
Synthetic
Steps
Expert
Epochs
Max Start
Epoch
(T+)
Synthetic
Batch
Size
Learning
Rate
Learning
Rate
(Step Size)
Starting
Synthetic
Step Size
Num
Eval
Eval
Iteration
$\lambda$
CIFAR-10 ConvNetD3 1 Y 50 2 2 - $10^{2}$ $10^{-7}$ $10^{-2}$ 5 100 0.0005
10 Y 30 2 20 - $10^{2}$ $10^{-4}$ $10^{-2}$ 5 100 0.02
50 N 30 2 40 - $10^{3}$ $10^{-5}$ $10^{-3}$ 5 100 0.0002
CIFAR-100 ConvNetD3 1 Y 20 3 20 - $10^{3}$ $10^{-5}$ $10^{-2}$ 5 100 0.001
10 N 20 2 20 - $10^{3}$ $10^{-5}$ $10^{-2}$ 5 100 0.02
50 Y 80 2 40 125 $10^{3}$ $10^{-5}$ $10^{-2}$ 5 100 0.002
Tiny ImageNet ConvNetD4 1 N 10 2 10 - $10^{4}$ $10^{-4}$ $10^{-2}$ 5 100 0.005
10 N 20 2 40 200 $10^{4}$ $10^{-4}$ $10^{-2}$ 3 200 0.02
50 N 20 2 40 300 $10^{4}$ $10^{-4}$ $10^{-2}$ 3 200 0.02
Table 6: Optimal hyperparameters for TESLA. A synthesis batch size of “-” means that we used the full support set at each synthesis step.
Dataset IPC
Matching
Steps
Teacher
Epochs
Max Start
Epoch
Synthetic
Batch
Size
Learning Rate
Learning
Rate
(Step Size)
Starting
Synthetic
Step Size
ZCA $\lambda$
CIFAR-10 1 50 2 3 - $10^{2}$ $10^{-7}$ $10^{-2}$ Y 0.01
10 30 2 20 - $10^{2}$ $10^{-4}$ $10^{-2}$ Y 0.002
50 26 3 40 - $10^{3}$ $10^{-5}$ $10^{-3}$ N 0.02
CIFAR-100 1 20 3 20 - $10^{3}$ $10^{-5}$ $10^{-2}$ Y 0.001
10 13 3 30 - $10^{3}$ $10^{-5}$ $10^{-2}$ N 0.002
50 50 2 40 100 $10^{3}$ $10^{-5}$ $10^{-2}$ Y 0.0002
Table 7: Optimal hyperparameters for FTD. A synthesis batch size of ‘-’ means that we used the full support set at each synthesis step.
Dataset IPC
Synthetic
Step
Expert
Epoch
Max Start
Epoch
Synthetic
Batch
Size
Learning
Rate
Learning
Rate
(Step Size)
Learning
Rate
Balance
coefficient
EMA
Decay
$\lambda$
CIFAR-10 1 50 2 2 - 100 $10^{-7}$ 0.01 0.3 0.9999 0.002
10 30 2 20 - 100 $10^{-5}$ 0.001 0.3 0.9995 0.002
50 30 2 40 - 1000 $10^{-5}$ 0.001 1 0.999 0.0002
CIFAR-100 1 40 3 20 - 1000 $10^{-5}$ 0.01 1 0.9995 0.002
10 20 2 40 - 1000 $10^{-5}$ 0.01 1 0.9995 0.0002
50 80 2 40 1000 1000 $10^{-5}$ 0.01 1 0.999 0.002
Table 8: Optimal hyperparameters for DATM.
Dataset Model IPC
Synthetic
Step
Expert
Epoch
Min Start
Epoch
Current
Max Start
Epoch
Max Start
Epoch
Synthetic
Batch Size
Learning
Rate
Learning
Rate
Num
Eval
Eval
Iteration
$\lambda$
CIFAR-10 ConvNetD3 1 80 2 0 4 4 10 5 100 5 500 0.0002
10 80 2 0 10 20 100 2 100 5 500 0.0005
50 80 2 0 20 40 500 2 1000 5 500 0.002
CIFAR-100 ConvNetD3 1 40 3 0 10 20 100 10 1000 5 500 0.02
Tiny ImageNet ConvNetD4 1 60 2 0 15 20 200 10 10000 5 500 0.002
10 60 2 10 50 50 250 10 100 3 500 0.002
50 80 2 40 70 70 250 10 100 3 500 0.002
B.2 Limitation
Computational Cost: Similar to other methods, we have not yet addressed the large computational cost associated with the dataset distillation. Our experiments were conducted on a mix of RTX 2080 Ti, RTX 3090, RTX 4090, NVIDIA A100, and NVIDIA V100 GPUs. The cost in terms of computational resources and time remains significant for large datasets and high IPC experiments. For example, distilling Tiny ImageNet using DATM with IPC = 1 requires approximately 150GB of GPU memory, and for IPC = 50, a single experiment can take nearly 24 hours to complete.
Hyperparameter Tuning: The selection of the $\lambda$ requires manual adjustment, which may involve additional costs. The extensive training durations and substantial GPU memory requirements make it challenging to conduct exhaustive experiments with multiple $\lambda$ values to identify the global optimum, given our computational resource limitations. By exploring a wider range of $\lambda$ values, it is possible to obtain better results.
B.3 Pseudocodes of adding SDC on Matching-based Distillation Methods
We provide detailed pseudocodes for GM-based methods and TM-based methods. We take DC as the standard GM-based method, and MTT as the standard TM-based method. The detailed pseudocodes are shown in Algorithm 1 for GM-based methods and Algorithm 2 for TM-based methods.
Algorithm 1 Gradient Matching with Sample Difficulty Correction
1:Training set $\mathcal{D}_{\mathsf{real}}$, category set $C$, classification cross-entropy loss $\mathcal{L}$, probability distribution for weights $P_{\theta}$, distance metric $\mathbf{D}$, regularization coefficient $\lambda$, number of steps $T$, learning rate $\eta$ for network parameters.
2:Initialize distilled data $\mathcal{D}_{\mathsf{syn}}\sim\mathcal{D}_{\mathsf{real}}$.
3:for each distillation step… do
4:    $\triangleright$ Initialize network $\theta_{0}\sim P_{\theta}$
5:    for $t=0\to T$ do
6:         for $c=0\to C-1$ do
7:             $\triangleright$ Sample a mini-batch of distilled images: $\mathcal{B}_{\mathsf{real}}^{c}\sim\mathcal{D}^{c}_{\mathsf{real}}$
8:             $\triangleright$ Sample a mini-batch of original images: $\mathcal{B}_{\mathsf{syn}}^{c}\sim\mathcal{D}^{c}_{\mathsf{syn}}$
9:             $\triangleright$ Compute $\mathcal{L}_{\mathcal{B}_{\mathsf{syn}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{\mathsf{real}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right]$, $\mathcal{L}_{\mathcal{B}_{\mathsf{real}}^{c}}=\mathbb{E}_{(x,y)\in\tilde{\mathcal{B}}_{\mathsf{real}}^{c}}\left[\mathcal{L}(x,y;\theta_{t})\right]$
10:             $\triangleright$ Compute gradient matching loss $L=\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{\mathsf{real}}^{c}},\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{\mathsf{syn}}^{c}}\right)+\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{B}_{\mathsf{syn}}}\right\|_{2}^{2}$
11:             $\triangleright$ Update $\mathcal{D}_{\mathsf{syn}}$ w.r.t. $L$
12:         end for
13:         $\triangleright$ Update network w.r.t. classification loss: $\theta_{t+1}=\theta_{t}-\eta\nabla\mathcal{L}_{\mathcal{D}_{\mathsf{syn}}}(\theta_{t})$
14:    end for
15:end for
16:distilled data $\mathcal{D}_{\mathsf{syn}}$
Algorithm 2 Trajectory Matching with Sample Difficulty Correction
1:Set of expert parameter trajectories trained on $\mathcal{D}_{\mathsf{real}}$$\{\tau_{i}^{*}\}$, the number of updates between starting and target expert params $M$, the number of updates to student network per distillation step $N$, differentiable augmentation function $\mathcal{A}$, maximum start epoch $T^{+}<T$, learning rate $\eta$ for network parameters, regularization coefficient $\lambda$, classification cross-entropy loss $\mathcal{L}$.
2:Initialize distilled data $\mathcal{D}_{\mathsf{syn}}\sim\mathcal{D}_{\mathsf{real}}$.
3:for each distillation step… do
4:    $\triangleright$ Sample expert trajectory: $\tau^{*}\sim\{\tau^{*}_{i}\}$ with $\tau^{*}=\{\theta_{t}^{\mathcal{D}_{\mathsf{real}}}\}_{0}^{T}$
5:    $\triangleright$ Choose random start epoch, $t\leq T^{+}$
6:    $\triangleright$ Initialize student network with expert params: $\theta_{t}^{\mathcal{D}_{\mathsf{syn}}}:=\theta_{t}^{\mathcal{D}_{\mathsf{real}}}$
7:    for $n=0\to N-1$ do
8:         $\triangleright$ Sample a mini-batch of distilled images: $\mathcal{B}_{\mathsf{syn}}\sim\mathcal{D}_{\mathsf{syn}}$
9:         $\triangleright$ Update student network w.r.t. classification loss: $\theta_{t+n+1}^{\mathcal{D}_{\mathsf{syn}}}=\theta_{t+n}^{\mathcal{D}_{\mathsf{syn}}}-\eta\nabla\mathcal{L}_{\mathcal{A}(\mathcal{B}_{\mathsf{syn}})}(\theta_{t+n}^{\mathcal{D}_{\mathsf{syn}}})$
10:    end for
11:    $\triangleright$ Compute loss between ending student and expert params: $L=\frac{\|\theta_{t+N}^{\mathcal{D}_{\mathsf{syn}}}-\theta_{t+M}^{\mathcal{D}_{\mathsf{real}}}\|_{2}^{2}}{\|\theta_{t+M}^{\mathcal{D}_{\mathsf{real}}}-\theta_{t}^{\mathcal{D}_{\mathsf{real}}}\|_{2}^{2}}+\lambda\left\|\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{\mathsf{syn}}}\right\|_{2}^{2}$
12:    $\triangleright$ Update $\mathcal{D}_{\mathsf{syn}}$ w.r.t. $L$
13:end for
14:distilled data $\mathcal{D}_{\mathsf{syn}}$
Appendix C Exploring the Effectiveness of SDC in Additional Experiments
C.1 More Results on the Cross-architecture Evaluation
To evaluate the performance of distilled datasets on different network architectures using SDC (marked as +SDC in the tables) and other methods (DATM and DSAC), we conducted cross-architecture evaluation experiments. We compared the effects of DATM and SDC on CIFAR-10, CIFAR-100, and Tiny ImageNet datasets, and the effects of DSAC and SDC on MNIST, FashionMNIST, and SVHN datasets. Finally, we further evaluated the performance differences between DSAC and SDC methods on MNIST, FashionMNIST, and SVHN datasets with IPC = 50. The cross-architecture evaluation experiments for DSAC and DATM, as well as the use of the SDC method on datasets with IPC = 1 of DATM and IPC = 10 of DSAC, can be found in Table 3.
The results of evaluating distilled datasets learned through DATM and SDC methods on CIFAR-10, CIFAR-100, and Tiny ImageNet datasets using ResNet-18, VGG-11, AlexNet, LeNet, and MLP networks are presented in Table 9. For instance, on the CIFAR-100 dataset, the accuracy of the VGG-11 network improved by 1.46%. It can be observed that the performance after applying SDC is generally better than DATM.
In our evaluation of distilled datasets learned through DSAC and SDC methods on MNIST, FashionMNIST, and SVHN datasets using the same network architectures, as detailed in Table 10, the results show that performance after applying SDC is superior to the DSAC method across datasets and network architectures. For example, on the MNIST dataset, the accuracy of the VGG-11 network improved by 1.37%, and on the SVHN dataset, the accuracy of the ResNet-18 improved by 0.61%.
Additionally, similar evaluation results on MNIST, FashionMNIST, and SVHN datasets with an IPC value of 50 are summarized in Table 11. For example, on the SVHN dataset, the accuracy of the LeNet network improved by 1.0%. It can be seen that with an increase in IPC value, the performance after applying SDC remains better in most cases, further demonstrating the superiority of the SDC method in dataset distillation.
Table 9: Cross-architecture evaluation. We evaluated distilled datasets with IPC = 10 learned through DATM w/ and w/o SDC on different networks.
Dataset Method ResNet-18 VGG-11 AlexNet LeNet MLP
CIFAR-10 DATM 36.48 37.32 33.19 32.56 27.21
+SDC 38.33 38.22 34.56 33.17 27.62
CIFAR-100 DATM 17.87 14.71 15.09 11.76 11.52
+SDC 18.97 16.17 15.73 12.44 11.87
Tiny ImageNet DATM 6.33 8.67 6.18 3.65 3.34
+SDC 7.20 9.13 6.89 3.88 3.40
Table 10: Cross-architecture evaluation. We evaluated distilled datasets with IPC = 1 learned through DSAC w/ and w/o SDC on different networks.
Dataset Method ResNet-18 VGG-11 AlexNet LeNet MLP
MNIST DSAC 88.58 79.58 83.63 83.46 72.78
+SDC 88.70 80.95 83.92 83.66 73.51
FashionMNIST DSAC 71.60 68.03 66.03 67.09 63.85
+SDC 71.70 68.82 66.47 67.19 64.93
SVHN DSAC 33.04 32.32 14.63 20.89 13.32
+SDC 33.65 33.84 17.18 22.40 13.86
Table 11: Cross-architecture evaluation. We evaluated distilled datasets with IPC = 50 learned through DSAC w/ and w/o SDC on different networks.
Dataset Method ResNet-18 VGG-11 AlexNet LeNet MLP
MNIST DSAC 97.97 98.53 97.95 97.58 94.70
+SDC 97.95 98.57 97.97 97.62 94.75
FashionMNIST DSAC 86.92 87.03 85.61 84.96 83.56
+SDC 86.96 87.20 85.58 85.36 83.78
SVHN DSAC 86.10 85.62 83.47 77.92 62.68
+SDC 86.32 85.86 83.85 78.92 63.47
C.2 More Results on the Adaptive Sample Difficulty Correction
Figure 13: The application of ASDC on (a) FTD and (b) DSA. Gradually increasing $\lambda$ gives better results than the baseline and the method after applying SDC.
The dynamic adjustment of SDC, when applied to both DSA and FTD, consistently outperforms both the baseline methods and the baseline methods with SDC applied. As shown in Figure 13, we logarithmically increased the $\lambda$ coefficient for DSA from 0.0002 to 0.002 over 1000 steps and for FTD from 0.002 to 0.008 over 10,000 iterations. The results clearly demonstrate that ASDC yields superior performance. Flexibly adjusting the sample difficulty correction by adaptively increasing $\lambda$ yields higher accuracy compared to the standard SDC and baseline methods.
C.3 Sensitivity Analysis of SDC coefficient $\lambda$.
In this section, we conducted extensive experiments to study the sensitivity of the hyperparameter $\lambda$. Specifically, we conducted experiments of DSA on SVHN dataset with IPC = 1, and DC on SVHN dataset with IPC = 10. As shown in Figure 14, the choice of $\lambda$ is not sensitive among different matching-based dataset distillation methods.
Figure 14: Sensitivity analysis of SDC coefficient $\lambda$ on different distillation methods. We evaluated the sensitivity across different $\lambda$s, and showed that the choice of $\lambda$ did not severely affect the final test performance. (a) DSA on SVHN dataset with IPC = 1 (b) DC on SVHN dataset with IPC = 10.
Appendix D Analytical Theory for Dataset Distillation
In this section, we introduce a theory, adapted from data pruning [42], to the context of dataset distillation within an expert-student perceptron framework, utilizing the tools of statistical mechanics. We investigate the challenge of classifying a dataset $\mathcal{D}_{\mathsf{real}}$ consisting of $d_{\mathsf{real}}$ samples $\{x_{i},y_{i}\}_{i=1,\ldots,d_{\mathsf{real}}}$, where the inputs $x_{i}\sim\mathcal{N}(0,I_{d})$ are i.i.d. zero-mean, unit-variance random Gaussian variables, and the labels $y_{i}=\text{sign}(\theta^{\mathcal{D}_{\mathsf{real}}^{\top}}x_{i})$ are generated by an expert perceptron $\theta^{\mathcal{D}_{\mathsf{real}}}\in\mathbb{R}^{d}$. We assume that the expert perceptron $\theta^{\mathcal{D}_{\mathsf{real}}}$ is randomly drawn from a uniform distribution on the sphere $\theta^{\mathcal{D}_{\mathsf{real}}}\sim\text{Unif}(\mathbb{S}^{d-1}(\sqrt{d}))$. Our analysis is situated within the high-dimensional statistics limit where $d,d_{\mathsf{real}}\to\infty$ but the ratio $\alpha_{\mathsf{real}}=d_{\mathsf{real}}/d$ remains $O(1)$.
Specifically, consider synthesizing a dataset by matching only the samples with the smallest margin $|z_{i}|=|\theta^{\mathsf{probe}^{\top}}x_{i}|$ along a probe student $\theta^{\mathsf{probe}}$. The distilled dataset will then follow a distribution $p(z)$ in the direction of $\theta^{\mathsf{probe}}$ while remaining isotropic in the null space of $\theta^{\mathsf{probe}}$. We assume, without loss of generality, that $\theta^{\mathsf{probe}}$ has developed some overlap with the expert, quantified by the angle $\gamma=\cos^{-1}\left(\frac{\theta^{\mathsf{probe}^{\top}}\theta^{\mathcal{D}_{\mathsf{real}}}}{\|\theta^{\mathsf{probe}}\|_{2}\|\theta^{\mathcal{D}_{\mathsf{real}}}\|_{2}}\right)$.
Once the dataset has been distilled, we consider training a new student $\theta^{\mathcal{D}_{\mathsf{syn}}}$ from scratch on this distilled dataset. A typical training algorithm aims to find the solution $\theta^{\mathcal{D}_{\mathsf{syn}}}$ which classifies the training data with maximal margin $\kappa=\min_{i}(\theta^{\mathcal{D}_{\mathsf{syn}}^{\top}}y_{i}x_{i})$. Our goal is to compute the generalization error $\varepsilon$ of this student, governed by the overlap between the student and the expert: $\varepsilon=\cos^{-1}(R)/\pi$, where $R=\frac{\theta^{\mathcal{D}_{\mathsf{syn}}^{\top}}\theta^{\mathcal{D}_{\mathsf{real}}}}{\|\theta^{\mathcal{D}_{\mathsf{syn}}}\|_{2}\|\theta^{\mathcal{D}_{\mathsf{real}}}\|_{2}}$.
We provide saddle point equations for the cosine similarity $R$ between the probe $\theta^{\mathcal{D}_{\mathsf{probe}}}$ and the expert $\theta^{\mathcal{D}_{\mathsf{real}}}$, which will be discussed in Section D.1 and Section D.2. For our simulations, we set the parameter dimension $d=200$ for perfect probe settings, and set $d=50$ for imperfect probe settings. We averaged 100 simulation results to verify the theory.
D.1 Perfect Expert-Teacher Settings
The solution is given by the following saddle point equations for perfect exeprt-teacher settings, i.e., $\gamma=0$. For any given $\alpha_{\mathsf{syn}}$, these equations can be solved for the order parameters $R,\kappa$. From these parameters, the generalization error can be computed as $\varepsilon=\cos^{-1}(R)/\pi$.
$R=\frac{2\alpha_{\mathsf{syn}}}{f\sqrt{2\pi}\sqrt{1-R^{2}}}\int_{-\infty}^{\kappa}Dt\ \exp\bigg{(}-\frac{R^{2}t^{2}}{2(1-R^{2})}\bigg{)}\bigg{[}1-\exp\bigg{(}-\frac{\gamma(\gamma-2Rt)}{2(1-R^{2})}\bigg{)}\bigg{]}(\kappa-t)$ (13)
$1-R^{2}=\frac{2\alpha_{\mathsf{syn}}}{f}\int_{-\infty}^{\kappa}Dt\ \bigg{[}H\bigg{(}-\frac{Rt}{\sqrt{1-R^{2}}}\bigg{)}-H\bigg{(}-\frac{Rt-\gamma}{\sqrt{1-R^{2}}}\bigg{)}\bigg{]}(\kappa-t)^{2}$ (14)
[…]
D.2 Imperfect Expert-Teacher Settings
We have shown the perfect student settings in Section D.1. When the probe student does not exactly match the expert, an additional parameter $\theta$ characterizes the angle between the probe student and the expert. Furthermore, an additional order parameter $\rho=\theta^{\mathcal{D}_{\mathsf{real}}^{\top}}\theta^{\mathcal{D}_{\mathsf{syn}}}$ represents the typical student-probe overlap, which must be optimized. Consequently, we derive three saddle point equations.
[…]
$\quad\quad\quad+\frac{1}{2\pi\Lambda}\exp\left(-\frac{\Delta(t,z)}{2\Lambda^{2}}\right)\bigg{(}\frac{\rho R-\cos\gamma}{1-\rho^{2}}\bigg{)}(\kappa-t)\bigg{>}_{z}$
Where,
$\Lambda$ $=\sqrt{\sin^{2}\gamma-R^{2}-\rho^{2}+2\rho R\cos\gamma},$ (15)
$\Gamma(t,z)$ $=z(\rho R-\cos\gamma)-t(R-\rho\cos\gamma),$
$\Delta(t,z)$ $=z^{2}\left(\rho^{2}+\cos^{2}\gamma-2\rho R\cos\gamma\right)+2tz(R\cos\gamma-\rho)+t^{2}\sin^{2}\gamma.$
The notation $\langle\cdot\rangle_{z}$ denotes an average over the pruned data distribution $p(z)$ for the probe student. For any given $\alpha_{\mathsf{syn}},p(z),\gamma$, these equations can be solved for the order parameters $R,\rho,\kappa$. From these parameters, the generalization error can be readily obtained as $\varepsilon=\cos^{-1}(R)/\pi$. Our simulation results are shown in Figure 15. Please refer [42] for detailed deductions.
Figure 15: Test error $\varepsilon$ as a function of the synthetic samples per parameter $\alpha_{\mathsf{syn}}$ and fraction of data synthesized $f$ in (a) the perfect expert setting ($\gamma=0$) (b) the perfect expert setting ($\gamma=10^{\circ}$) (c) the perfect expert setting ($\gamma=20^{\circ}$).
Appendix E Visualization Results
Additionally, we show our visualization of distilled datasets by adding SDC into current matching-based methods, as shown in Figure 16, Figure 17, Figure 18, Figure 19, and Figure 20.
Figure 16: (FTD + SDC, CIFAR-10, IPC = 10) Visualization of distilled images.
Figure 17: (DATM + SDC, Tiny ImageNet, IPC = 1, 1 / 2) Visualization of distilled images.
Figure 18: (DATM + SDC, Tiny ImageNet, IPC = 1, 2 / 2) Visualization of distilled images.
Figure 19: (DSA + SDC, SVHN, IPC = 10) Visualization of distilled images.
Figure 20: (DC + SDC, FashionMNIST, IPC = 10) Visualization of distilled images.